# appendixReferences

### v3

@@ 3.3 Point Estimation
Consider an experiment in which $100$ units are divided equally between treatment and control, where the average outcome for the treated and control groups are respectively $0.8$ (40/50) and $0.4$ (20/50). Units who are unaffected by treatment have fixed outcomes, and are equally likely to be assigned to treatment or control. It follows that up to sampling variation, such units should be evenly distributed between treatment and control even after stratifying by outcome. Therefore, out of the 40 treated units whose outcome equals 1, at most roughly 20 such units could be unaffected by treatment, if they are to be equally represented in the control group. Likewise out of the 30 control units whose outcome equals 0, at most roughly 10 could be unaffected by treatment, if they are to be equally represented in the treatment group. This yields an upper bound of $(20+10)\times 2=60$ units who are unaffected by treatment, or equivalently a lower bound of 40 units who are affected by treatment. The lower bound corresponds to the difference in means of 0.4 between treatment and control. The limiting quantities that were used to construct this bound, namely the 20 control units with outcome equal to 1, and the 10 treated units with outcome equal to 0, are equivalent to the set $\{i:Y_{i}\neq X_{i}\}$ utilized by the estimator $\hat{\Phi}_{2}$.
Continuing the example, we describe a generative model in which the contrast in outcomes is tight for the number of units affected by treatment. In this model, the effect on unit $i$ of $X_{i}$, the contrast variable appearing in (12), is monotone–that is, not positive for some units and negative for others. We then describe two models under which the lower bound is loose, and which suggest that non-vaccuous upper bounds are not identifiable.
  1. 1.
@@
    (Nonnegative effects) Of the 100 units in the experiment, let 40 units have outcome 1 regardless of treatment, 20 units have outcome 0 regardless of treatment, and the remaining 40 units have outcome 1 if treated and 0 if control. Under this model, the expected outcomes for treated and control can be seen to be $0.8$ and $0.4$ as described above, and the empirical diff-in-means will equal (up to sampling variation) the fraction of units affected by treatment.
  2. 2.
@@
    (Cancellation of effects) Of the 100 units in the experiment, let 20 units have outcome 1 regardless of treatment, 20 units have outcome 0 if treated and 1 if control, and the remaining 60 units have outcome 1 if treated and 0 if control. Under this model, direct treatment affects some units positively and others negatively. While the expected outcomes match the previous example, here the diff-in-means is a loose lower bound on the fraction of units affected by treatment, which is 0.8.
  3. 3.
    (Pathological interference) Let $Y_{i}$ equal
    $Y_{i}$ =$=\begin{cases}\alpha_{i1}X_{i}+\alpha_{i0}(1-X_{i})&\textup{ if $\sum_{i}X_{i}$ is even}\\ \beta_{i1}X_{i}+\beta_{i0}(1-X_{i})&\textup{ if $\sum_{i}X_{i}$ is odd}\end{cases}$ (13)
@@
    where $(\alpha_{i1},\alpha_{i0})$ describe the potential outcomes of the first example, equaling $(1,1)$ for units 1-40, (0,0) for units 41-60, and $(1,0)$ for units 61-100, while $(\beta_{i1},\beta_{i0})$ takes values $(1,0)$ and $(0,1)$ in some arbitrary proportion. If half of the units are assigned to treatment under the experiment design, then the outcomes will be indistinguishable from those of the first example; however, unlike the first example, all 100 units are affected by treatment.
Using Network Information
@@
To utilize $W$, the thresholded number of treated neighbors given by (2), we may let $\hat{\Phi}_{1}$ and $\hat{\Phi}_{2}$ equal
$\hat{\Phi}_{k}=\sum_{i:\phi_{i}^{*}=1}\frac{1\{X_{i}=W_{i}=Y_{ki}\}}{{P}(X_{i}=W_{i}=Y_{ki})},\qquad k=1,2.$ (14)
For this version of $\hat{\Phi}_{k}$, we may show that $\hat{\tau}$ remains a consistent lower bound on $\tau^{\textup{basic}}$ under the conditions of Theorem 2, but converges to a different contrast given by
$\hat{\tau}=\left|\sum_{i=1}^{N}\frac{X_{i}W_{i}}{P(X_{i}W_{i}=1)}Y_{i}-\sum_{i=1}^{N}\frac{(1-X_{i})(1-W_{i})}{P(X_{i}=0,W_{i}=0)}Y_{i}\right|+o_{P}(N)$ (15)
This is the IPW difference in outcomes between units whose exposure variable $(X_{i},W_{i})$ equals (1,1) versus (0,0). If the effects of direct and indirect treatment reinforce each other, this contrast may be larger than the comparison of treated and control given by (12), implying a larger and tighter lower bound on $\tau^{\textup{basic}}$.
To give a single formula that includes both (5) and (14) as special cases, let $\hat{\Phi}_{k}$ equal
$\hat{\Phi}_{k}=\sum_{i:\phi_{i}^{*}=1}\frac{1\{Z_{i}=Y_{ki}\}}{P(Z_{i}=Y_{ki})},$ (16)
where $Z_{i}\in\{1,0,-1\}$ divides the units into 3 groups, as a function of each unit’s direct and indirect treatment. Under (16), $\hat{\tau}$ will converge to an IPW contrast between the units for which $Z_{i}$ equals 1 and 0 (“effective treatment” versus “effective control”), while units for which $Z_{i}=-1$ are not used. Letting $Z_{i}=X_{i}$ causes (16) to be equivalent to (5), while (14) is recovered by letting $Z_{i}$ equal
$Z_{i}=\begin{cases}1&\textup{ if $X_{i}=W_{i}=1$}\\ 0&\textup{ if $X_{i}=W_{i}=0$}\\ -1&\textup{ if $X_{i}\neq W_{i}$}\end{cases}$
@@ 3.4 Hájek Normalization
To reduce the variance of our IPW-based estimators, we may use Hájek normalization (Aronow and Samii, , 2017; Gao and Ding, , 2023). To construct a Hájek-normalized version of $\hat{\Phi}_{k}$ given by (16), we first observe that it equals
$\hat{\Phi}_{k}=\sum_{i:\phi_{i}^{*}=1}\left(\frac{Z_{i}Y_{ki}}{P(Z_{i}=1)}+\frac{(1-Z_{i})(1-Y_{ki})}{P(Z_{i}=0)}\right)$ (17)
where we have divided the event $\{Z_{i}=Y_{ki}\}$ into sub-events $\{Z_{i}=Y_{ki}=1\}$ and $\{Z_{i}=Y_{ki}=0\}$, and replaced the propensity score $P(Z_{i}=Y_{ki})$ by $Y_{ki}P(Z_{i}=1)+(1-Y_{ki})P(Z_{i}=0)$, which can be done because $Y_{ki}$ is deterministic when $\phi_{i}^{*}=1$. Eq. (17) suggests a Hájek normalized version given by
$\hat{\Phi}_{k}^{{\textup{Haj}}}=\sum_{i:\phi_{i}^{*}=1}\left[\frac{N}{\hat{N}_{1}}\cdot\frac{Z_{i}}{P(Z_{i}=1)}\cdot Y_{ki}+\frac{N}{\hat{N}_{0}}\cdot\frac{1-Z_{i}}{P(Z_{i}=0)}\cdot(1-Y_{ki})\right],$ (18)
@@
To construct a Hájek-weighted point estimate for $\tau^{\textup{basic}}$, we can substitute $\hat{\Phi}_{k}^{{\textup{Haj}}}(\phi)$ for $\hat{\Phi}_{k}(\phi)$ in the expression for $\hat{\tau}$ given by (8). In Theorem 4, we show that the resulting $\hat{\tau}$ exactly equals a Hájek-weighted comparison of units, given by
$\hat{\tau}=\left|\sum_{i=1}^{N}\frac{N}{\hat{N}_{1}}\cdot\frac{Z_{i}}{P(Z_{i}=1)}Y_{i}-\sum_{i=1}^{N}\frac{N}{\hat{N}_{0}}\cdot\frac{(1-Z_{i})}{P(Z_{i}=0)}Y_{i}\right|.$
@@ 3.5 Inference and Variance Estimation
If $\hat{\Phi}_{1}^{\textup{Haj}}$ and $\hat{\Phi}_{2}^{\textup{Haj}}$ are asymptotically normal, with consistent variance estimators denoted by $V_{1}$ and $V_{2}$, then by combining 1-sided confidence intervals it holds with probability converging to at least $1-\alpha$ that
$\Phi$ $\leq\min\left\{\hat{\Phi}_{1}^{\textup{Haj}}+z_{1-\frac{\alpha}{2}}\sqrt{V_{1}},\,\hat{\Phi}_{2}^{\textup{Haj}}+z_{1-\frac{\alpha}{2}}\sqrt{V_{2}}\right\}.$ (19)
To find a one-sided confidence interval for $\tau^{\textup{basic}}=N-\Phi$, we will upper bound $\Phi$ by (19) maximized over hypotheses for the unknown $\phi^{*}$. This yields the confidence statement that with probability converging to at least $(1-\alpha)$, it holds that
$\tau^{\textup{basic}}\geq N-\min\Big{(}\max_{\phi\in\{0,1\}^{N}}\hat{\Phi}_{1}^{\textup{Haj}}(\phi)+z_{1-\frac{\alpha}{2}}\sqrt{V_{1}(\phi)},\,\max_{\phi\in\{0,1\}^{N}}\hat{\Phi}_{2}^{\textup{Haj}}(\phi)+z_{1-\frac{\alpha}{2}}\sqrt{V_{2}(\phi)}\,\Big{)}$ (20)
where $V_{k}(\phi)$ denotes $V_{k}$, evaluated under the hypothesis that $\phi^{*}=\phi$. We give an expression for $V_{k}$ next.
@@
The expression for $\hat{\Phi}_{k}^{\textup{Haj}}$ given by (18) is equivalent to
$\hat{\Phi}_{k}^{{\textup{Haj}}}=\sum_{i:\phi_{i}^{*}=1}\frac{N}{\hat{N}_{Y_{ki}}}\cdot\frac{1\{Z_{i}=Y_{ki}\}}{P(Z_{i}=Y_{ki})}$
It follows that the variance of $\hat{\Phi}_{k}$ can be written as a sum of covariance terms,
${\operatorname{Var}\,}\hat{\Phi}_{k}^{\textup{Haj}}=\sum_{i:\phi_{i}^{*}=1}\sum_{j:\phi_{j}^{*}=1}{\operatorname{Cov}}\left(\frac{N}{\hat{N}_{Y_{ki}}}\cdot\frac{1\{Z_{i}=Y_{ki}\}}{P(Z_{i}=Y_{ki})}\ ,\ \frac{N}{\hat{N}_{Y_{ki}}}\cdot\frac{1\{Z_{i}=Y_{ki}\}}{P(Z_{i}=Y_{ki})}\right)$ (21)
To construct the one-sided interval (20), we propose $V_{k}$ given by
$V_{k}=\sum_{i:\phi_{i}^{*}=1}\sum_{j:\phi_{j}^{*}=1}{\operatorname{Cov}}\left(\frac{N}{\hat{N}_{Y_{ki}}}\cdot\frac{1\{Z_{i}=Y_{ki}\}}{P(Z_{i}=Y_{ki})}\ ,\ \frac{N}{\hat{N}_{Y_{ki}}}\cdot\frac{1\{Z_{i}=Y_{ki}\}}{P(Z_{i}=Y_{ki})}\right)\\ \cdot\frac{1\{Z_{i}=Y_{ki},Z_{j}=Y_{kj}\}}{P(Z_{i}=Y_{ki},Z_{j}=Y_{kj})}$ (22)
Comparing (21) and (22), it can be seen that $V_{k}$ is a Horvitz-Thompson estimator which samples the covariance terms appearing in $\operatorname{Var}\hat{\Phi}_{k}$. Hence it is unbiased provided that the sampling probabilities $P(Z_{i}=Y_{ki},Z_{j}=Y_{kj})$ are non-zero for all $i,j\in[N]^{2}$. Should this condition not hold, our variance estimators will be conservatively biased, as will be established in Theorem 5.
To evaluate (22), we may use the following expressions for the covariance and propensity values, which hold whenever $Y_{ki}$ and $Y_{kj}$ are deterministic, and hence for all $(i,j)$ such that $\phi_{i}^{*}\phi_{j}^{*}=1$,
${\operatorname{Cov}}\left(\frac{N}{\hat{N}_{Y_{ki}}}\cdot\frac{1\{Z_{i}=Y_{ki}\}}{P(Z_{i}=Y_{ki})}\ ,\ \frac{N}{\hat{N}_{Y_{ki}}}\cdot\frac{1\{Z_{i}=Y_{ki}\}}{P(Z_{i}=Y_{ki})}\right)=\\ \sum_{a=0}^{1}\sum_{b=0}^{1}1\{Y_{ik}=a,Y_{jk}=b\}{\operatorname{Cov}}\left(\frac{N}{\hat{N}_{a}}\cdot\frac{1\{Z_{i}=a\}}{P(Z_{i}=a)}\ ,\ \frac{N}{\hat{N}_{b}}\cdot\frac{1\{Z_{i}=b\}}{P(Z_{i}=b)}\right)$ (23)
@@
$P(Z_{i}=Y_{ki},Z_{j}=Y_{kj})=\sum_{a=0}^{1}\sum_{b=0}^{1}1\{Y_{ik}=a,Y_{jk}=b\}P(Z_{i}=a,Z_{j}=b).$ (24)
Using these expressions, we can evaluate $V_{1}(\phi)$, which equals $V_{1}$ under the hypothesis that $\phi^{*}=\phi$, for any $\phi\in\{0,1\}^{N}$.
The motivation for our sampling-based variance estimator is that it takes small value when the treated and control units have large differences in average outcome. Analogous to the intuition portrayed in Figure 1, when the diff-in-means is large, the number of pairs $(i,j)$ such that the event $\{Z_{i}=Y_{ki},Z_{j}=Y_{kj}\}$ occurs must be small for either $k=1$ or $k=2$, and hence $V_{1}(\phi)$ or $V_{2}(\phi)$ will be small, even if maximized over $\phi\in\{0,1\}^{N}$.
We discuss some conceptual differences between the variance estimator of our setting and that of treatment effect estimation. Generally, the variance of a treatment effect estimator involves products between pairs of potential outcomes that cannot be simultaneously observed, such as the outcome of the same unit under both treatment and control. Such product terms cannot be estimated in an unbiased manner, but rather must be conservatively bounded (Harshaw et al., , 2021). These difficulties do not arise in the variance expression (21), as it involves only the outcomes of units who are unaffected by treatment. As a result, $V_{k}$ will be unbiased under conditions to be described in Theorem 5, but must itself be bounded over the unknown set of units who are unaffected by treatment, as is done in (20).

### v4

@@ Discussion
(SUTVA with nonnegative effects) Of the 100 units in the experiment, suppose that units 1-60 are unaffected by treatment, and units 61-100 are affected by their own treatment, according to
$f_{i}(X)=\begin{cases}1&1\leq i\leq 40\\ 0&41\leq i\leq 60\\ X_{i}&61\leq i\leq 100\end{cases}$
@@
Under this model, $40$ units are affected by treatment, and the expected outcomes for treated and control can be seen to be $0.8$ and $0.4$. Hence the diff-in-means will equal the fraction of units affected by treatment (up to sampling variation).
Example 2.
(SUTVA with cancellation of effects) Suppose that units 1-20 are unaffected by treatment, while units 21-40 are affected negatively by their own treatment, and units 41-100 are affected positively,
$f_{i}(X)=\begin{cases}1&1\leq i\leq 20\\ 1-X_{i}&21\leq i\leq 40\\ X_{i}&41\leq i\leq 100\end{cases}$
@@
While the expected outcomes match the previous example, equaling $0.8$ and $0.4$ for treated and control, the diff-in-means is a loose lower bound on the fraction of units affected by treatment, which is 0.8.
Example 3.
(Interference) Let the outcome models be given by
[…]
@@
where $(\alpha_{i1},\alpha_{i0})$ match the potential outcomes of the first example, equaling $(1,1)$ for units 1-40, (0,0) for units 41-60, and $(1,0)$ for units 61-100, while $(\beta_{i1},\beta_{i0})$ takes values $(1,0)$ and $(0,1)$ in some arbitrary proportion. If half of the units are randomly assigned to treatment, then $\sum_{i}X_{i}$ is even and the outcomes will be indistinguishable from those of the first example; however, unlike the first example, all 100 units are affected by treatment. In this example $\Delta$ does not concentrate at its expectation, but is nonetheless a valid lower bound for $\tau^{\textup{basic}}$.
3.4 Using Network Information
Given an elicited social network, we may consider contrasts other than $\Delta$, based on a new variable $Z=(Z_{1},\ldots,Z_{N})$ which encodes the direct and indirect treatment of each unit. We will consider $Z\in\{1,0,-1\}^{N}$, where $1$ and 0 respectively denote generalized versions of treatment and control, and 1-1 denotes that the unit will be excluded from comparison.
@@ 3.4 Using Network Information
For example, to use $W$, the thresholded number of treated neighbors given by (2), we can define $Z$ according to
$Z_{i}=\begin{cases}1&\textup{ if $X_{i}=W_{i}=1$}\\ 0&\textup{ if $X_{i}=W_{i}=0$}\\ -1&\textup{ if $X_{i}\neq W_{i}$}\end{cases}$ (9)
Under $Z$ so defined, unit $i$ receives “effective treatment” if $X_{i}=1$ and $W_{i}=1$; receives “effective control” if $X_{i}=0$ and $W_{i}=0$; and otherwise is excluded from comparison.
Given $Z$, we can redefine $H=(H_{1},\ldots,H_{N})$ to use $Z$ instead of $X$,
[…]
which induces new versions of $\hat{\tau}_{1}$ and $\hat{\tau}_{2}$ following (5) as before. Proposition 1 can then be adapted to show that the difference in outcomes between effective treatment and effective control, given by
$\Delta^{Z}=\sum_{i=1}^{N}\left(\frac{1\{Z_{i}=1\}}{P(Z_{i}=1)}-\frac{1\{Z_{i}=0\}}{P(Z_{i}=0)}\right)Y_{i}$
is an asymptotic lower bound on the maximum of $\hat{\tau}_{1}$ and $\hat{\tau}_{2}$.
Proposition 2.
Let $Z$ be given by (9) and let the total weights of effective treatment and control converge to their expectations,
$\sum_{i=1}^{N}\frac{1\{Z_{i}=1\}}{P(Z_{i}=1)}=N+O_{P}(N^{1/2})\quad\textup{ and }\quad\sum_{i=1}^{N}\frac{1\{Z_{i}=0\}}{P(Z_{i}=0)}=N+O_{P}(N^{1/2})$
and let $\hat{\tau}_{1}$ and $\hat{\tau}_{2}$ be given by (5) for $H$ given by (10). Then it holds that
$\left|\sum_{i=1}^{N}\left(\frac{1\{Z_{i}=1\}}{P(Z_{i}=1)}-\frac{1\{Z_{i}=0\}}{P(Z_{i}=0)}\right)Y_{i}\right|\leq\max(\hat{\tau}_{1},\hat{\tau}_{2})+O_{P}(N^{1/2})$ (11)
If $Z$ is a better predictor of the unit outcomes than $X$, then $\Delta^{Z}$ may be larger than the comparison of direct treatment and direct control given by $\Delta$. This results in a larger and tighter lower bound on $\tau^{\textup{basic}}$, assuming consistency of $\hat{\tau}_{1}$ and $\hat{\tau}_{2}$.
@@ 3.5 Hájek Normalization
To reduce the variance of our IPW-based estimators, we may use Hájek normalization (Aronow and Samii, , 2017; Gao and Ding, , 2023). As a preliminary step, let $Y_{ki}$ denote transformed outcomes given by
$Y_{ki}=\begin{cases}Y_{i}&\textup{ if }k=1\\ 1-Y_{i}&\textup{ if }k=2,\end{cases}\qquad i\in[N],\ k=1,2,$ (12)
This allows $\hat{\tau}_{1}$ and $\hat{\tau}_{2}$ using $H$ given by (10) to be written with a single formula
$\hat{\tau}_{k}=N-\sum_{i\in{\mathcal{I}}}\frac{1\{Z_{i}=Y_{ki}\}}{P(Z_{i}=Y_{ki})},\qquad k=1,2,$ (13)
which will simplify our discussion of Hájek normalization for $\hat{\tau}_{1}$ and $\hat{\tau}_{2}$.
To construct a Hájek-normalized version of $\hat{\tau}_{k}$ as given by (13), we first observe that it equals
$\hat{\tau}_{k}=N-\sum_{i\in{\mathcal{I}}}\left(\frac{1\{Z_{i}=1,Y_{ki}=1\}}{P(Z_{i}=1)}+\frac{1\{Z_{i}=0,Y_{ki}=0\}}{P(Z_{i}=0)}\right)$ (14)
where we have divided the event $\{Z_{i}=Y_{ki}\}$ into sub-events $\{Z_{i}=Y_{ki}=1\}$ and $\{Z_{i}=Y_{ki}=0\}$, and replaced the propensity score $P(Z_{i}=Y_{ki})$ by
$P(Z_{i}=Y_{ki})=Y_{ki}P(Z_{i}=1)+(1-Y_{ki})P(Z_{i}=0),\qquad\forall\ i\in{\mathcal{I}},$
which can be done because $Y_{ki}$ is constant in $X$ for $i\in{\mathcal{I}}$. Eq. (14) suggests a Hájek normalized version given by
$\hat{\tau}_{k}^{{\textup{Haj}}}=N-\sum_{i\in{\mathcal{I}}}\left[\frac{N}{\hat{N}_{1}}\frac{1\{Z_{i}=1\}}{P(Z_{i}=1)}\cdot Y_{ki}+\frac{N}{\hat{N}_{0}}\frac{1\{Z_{i}=0\}}{P(Z_{i}=0)}\cdot(1-Y_{ki})\right],$ (15)
@@
In Theorem 1, we show that $\max(\hat{\tau}_{1}^{\textup{Haj}},\hat{\tau}_{2}^{\textup{Haj}})$ is lower bounded by a Hájek-weighted comparison of units,
$\left|\sum_{i=1}^{N}\left(\frac{N}{\hat{N}_{1}}\frac{Z_{i}}{P(Z_{i}=1)}-\frac{N}{\hat{N}_{0}}\frac{(1-Z_{i})}{P(Z_{i}=0)}\right)Y_{i}\right|\leq\max(\hat{\tau}_{1}^{\textup{Haj}}\,,\,\hat{\tau}_{2}^{\textup{Haj}})$
noting that the inequality holds non-asymptotically, in contrast to (7) and (11).
@@ 3.6 Inference and Variance Estimation
If $\hat{\tau}_{1}^{\textup{Haj}}$ and $\hat{\tau}_{2}^{\textup{Haj}}$ are asymptotically normal, with consistent variance estimators denoted by $\hat{V}_{1}$ and $\hat{V}_{2}$, then by combining 1-sided confidence intervals it holds with probability converging to at least $1-\alpha$ that
$\tau^{\textup{basic}}$ $\geq\max\left\{\hat{\tau}_{1}^{\textup{Haj}}-z_{1-\frac{\alpha}{2}}\sqrt{\hat{V}_{1}},\,\hat{\tau}_{2}^{\textup{Haj}}-z_{1-\frac{\alpha}{2}}\sqrt{\hat{V}_{2}}\right\}.$ (16)
As the right hand side of (16) requires knowledge of $\mathcal{I}$}, it cannot be computed.
To construct a computable one-sided confidence interval for $\tau^{\textup{basic}}$, we will lower bound (16) by minimizing over all possible hypotheses for the unknown $\mathcal{I}$}. Doing so results in the confidence statement that with probability at least $1-\alpha$,
$\tau^{\textup{basic}}\geq\max\left\{\min_{\phi\in\{0,1\}^{N}}\hat{\tau}_{1}^{\textup{Haj}}(\phi)-z_{1-\frac{\alpha}{2}}\sqrt{\hat{V}_{1}(\phi)},\,\min_{\phi\in\{0,1\}^{N}}\hat{\tau}_{2}^{\textup{Haj}}(\phi)-z_{1-\frac{\alpha}{2}}\sqrt{\hat{V}_{2}(\phi)}\,\right\},$ (17)
where $\hat{\tau}_{k}^{\textup{Haj}}(\phi)$ and $\hat{V}_{k}(\phi)$ denote $\hat{\tau}_{k}^{\textup{Haj}}$ and $\hat{V}_{k}$ evaluated under the hypothesis that ${\mathcal{I}}=\{i:\phi_{i}=1\}$ for $\phi\in\{0,1\}^{N}$.
To complete our discussion of interval estimation, we next give an expression for $\hat{V}_{k}$, and then discuss computation.
@@
The expression for $\hat{\tau}_{k}^{\textup{Haj}}$ given by (15) is equivalent to
$\hat{\tau}_{k}^{{\textup{Haj}}}=N-\sum_{i\in{\mathcal{I}}}\frac{N}{\hat{N}_{Y_{ki}}}\cdot\frac{1\{Z_{i}=Y_{ki}\}}{P(Z_{i}=Y_{ki})}$
It follows that the variance of $\hat{\tau}_{k}^{\textup{Haj}}$ can be written as a sum of covariance terms,
${\operatorname{Var}\,}\hat{\tau}_{k}^{\textup{Haj}}=\sum_{i\in{\mathcal{I}}}\sum_{j\in{\mathcal{I}}}{\operatorname{Cov}}\left(\frac{N}{\hat{N}_{Y_{ki}}}\cdot\frac{1\{Z_{i}=Y_{ki}\}}{P(Z_{i}=Y_{ki})}\ ,\ \frac{N}{\hat{N}_{Y_{kj}}}\cdot\frac{1\{Z_{j}=Y_{kj}\}}{P(Z_{j}=Y_{kj})}\right)$ (18)
To construct the one-sided interval (17), we propose $\hat{V}_{k}$ given by
$\hat{V}_{k}=\sum_{i\in{\mathcal{I}}}\sum_{j\in{\mathcal{I}}}{\operatorname{Cov}}\left(\frac{N}{\hat{N}_{Y_{ki}}}\frac{1\{Z_{i}=Y_{ki}\}}{P(Z_{i}=Y_{ki})}\ ,\ \frac{N}{\hat{N}_{Y_{kj}}}\frac{1\{Z_{j}=Y_{kj}\}}{P(Z_{j}=Y_{kj})}\right)\cdot\frac{1\{Z_{i}=Y_{ki},Z_{j}=Y_{kj}\}}{P(Z_{i}=Y_{ki},Z_{j}=Y_{kj})}$ (19)
Comparing (18) and (19), it can be seen that $\hat{V}_{k}$ is a Horvitz-Thompson estimator which samples the covariance terms appearing in $\operatorname{Var}\hat{\tau}_{k}^{\textup{Haj}}$. Hence it is unbiased provided that the sampling probabilities $P(Z_{i}=Y_{ki},Z_{j}=Y_{kj})$ are non-zero for all $i,j\in{\mathcal{I}}$. Should this condition not hold due to dependencies between the entries of $Z$, our variance estimators will be conservatively biased, as will be established in Theorem 2.
We use the estimated variance $\hat{V}_{k}$, rather than the actual variance given by (18), in the 1-sided interval given by (17). This is because (17) involves a minimization over all possible hypotheses for the unknown set $\mathcal{I}$}, and $\hat{V}_{k}$ can be less sensitive than the actual variance to this optimization, resulting in a less conservative interval. This is particularly the case when the contrast $\Delta^{Z}$ is large, as then either $\hat{V}_{1}$ or $\hat{V}_{2}$ involves only a small number of sampled covariance terms.
To evaluate $\hat{V}_{k}$ given by (19), we may use the following expressions for the covariance and propensity values which hold for all $i,j\in{\mathcal{I}}$,
${\operatorname{Cov}}\left(\frac{N}{\hat{N}_{Y_{ki}}}\cdot\frac{1\{Z_{i}=Y_{ki}\}}{P(Z_{i}=Y_{ki})}\ ,\ \frac{N}{\hat{N}_{Y_{kj}}}\cdot\frac{1\{Z_{j}=Y_{kj}\}}{P(Z_{j}=Y_{kj})}\right)=\\ \sum_{a=0}^{1}\sum_{b=0}^{1}1\{Y_{ki}=a,Y_{kj}=b\}\ {\operatorname{Cov}}\left(\frac{N}{\hat{N}_{a}}\cdot\frac{1\{Z_{i}=a\}}{P(Z_{i}=a)}\ ,\ \frac{N}{\hat{N}_{b}}\cdot\frac{1\{Z_{j}=b\}}{P(Z_{j}=b)}\right)$ (20)
@@
$P(Z_{i}=Y_{ki},Z_{j}=Y_{kj})=\sum_{a=0}^{1}\sum_{b=0}^{1}1\{Y_{ki}=a,Y_{kj}=b\}\ P(Z_{i}=a,Z_{j}=b).$ (21)
The expressions (20) and (21) hold because $Y_{ki}$ and $Y_{kj}$ are constants when $i,j\in{\mathcal{I}}$.